# Synthetic Data for Robust Stroke Segmentation

### v2

@@ 4.5 Ablation: To mix or not to mix?
For all experiments shown thus far, the Synth model used a mix of both synthetic data and the real ATLAS dataset. In order to evaluate whether this decision is justified, an ablation is performed where the Synth model using mixed real/synthetic data is compared to a model trained with only synthetic data. This model is trained in the exact same manner as described previously for the baseline and Synth models. Results for the test datasets ATLAS, ISLES 2015 and ARC are shown in Tables 5, 6 and 7 respectively. In the in-domain ATLAS T1w data, the Synth model without mixed real data performs significantly worse than both the mixed data Synth model and the baseline model. In the ISLES 2015 dataset, the model without real data performs significantly better in Dice for the T1w modality. For all other modalities, there is no significant difference between the two Synth models in Dice score. Finally, in the ISLES dataset the model without any real data again performs significantly worse than both others in the near-domain T1w images. It does however perform significantly better than both others for the T2 data.
@@
Figure 6: Dice and 95th-percentile Hausdorff Distance for all reported datasets, for models trained on different combinations of real/synthetic data.
@@
Table 5: Median results for an ablation study on the ATLAS test set (N=131). Best score shown in bold. Student’s t distribution 95% confidence intervals given in brackets.
Modality Model Dice HD95
T1w Baseline 0.575 (0.523-0.627) 23.5 (13.9-33.2)
Synth (no real data) 0.197 (0.155-0.239) 63.7 (59.6-67.9)
Synth (mix real/synth data) 0.482 (0.431-0.534) 22.6 (12.0-33.3)
Table 6: Median results for an ablation study on the ISLES 2015 test set (N=28). Best score shown in bold. Student’s t distribution 95% confidence intervals given in brackets.
Modality Model Dice HD95
T1w Baseline 0.000 (0.000-0.074) 67.0 (39.7-94.2)
Synth (no real data) 0.304 (0.213-0.396) 58.6 (48.5-68.6)
Synth (mix real/synth data) 0.110 (0.012-0.208) 52.5 (29.9-75.0)
T2w Baseline 0.000 (0.000-0.005) 62.7 (56.3-69.1)
Synth (no real data) 0.074 (0.000-0.184) 72.3 (65.7-78.9)
Synth (mix real/synth data) 0.111 (0.007-0.216) 75.0 (68.4-81.7)
FLAIR Baseline 0.000 (0.000-0.000) 65.8 (56.9-74.7)
Synth (no real data) 0.372 (0.255-0.489) 56.1 (49.5-62.7)
Synth (mix real/synth data) 0.212 (0.085-0.340) 56.1 (38.7-73.4)
DWI Baseline 0.000 (0.000-0.000) 73.1 (57.5-88.7)
Synth (no real data) 0.082 (0.000-0.168) 83.7 (76.3-91.0)
Synth (mix real/synth data) 0.056 (0.000-0.144) 84.9 (77.2-92.6)
Ensemble Baseline 0.000 (0.000-0.000) 256.0 (237.2-256.0)
Synth (no real data) 0.272 (0.157-0.388) 60.1 (49.7-70.5)
Synth (mix real/synth data) 0.423 (0.302-0.545) 47.3 (24.2-70.4)
@@
Table 7: Median results for an ablation study on the ARC test set (N=229). Best score shown in bold. Student’s t distribution 95% confidence intervals given in brackets.
Modality Model Dice HD95
T1w Baseline 0.752 (0.715-0.790) 8.6 (2.3-14.9)
Synth (no real data) 0.467 (0.430-0.504) 57.1 (52.9-61.3)
Synth (mix real/synth data) 0.723 (0.684-0.762) 11.0 (4.0-18.0)
T2w Baseline 0.004 (0.000-0.014) 74.2 (71.1-77.2)
Synth (no real data) 0.626 (0.588-0.665) 46.1 (42.2-50.0)
Synth (mix real/synth data) 0.268 (0.233-0.302) 64.0 (60.5-67.4)
FLAIR Baseline 0.120 (0.076-0.163) 63.0 (54.1-71.9)
Synth (no real data) 0.096 (0.070-0.123) 57.6 (49.2-65.9)
Synth (mix real/synth data) 0.141 (0.097-0.185) 52.2 (43.3-61.1)
Ensemble Baseline 0.034 (0.013-0.054) 40.4 (33.5-47.3)
Synth (no real data) 0.597 (0.562-0.631) 44.0 (39.9-48.2)
Synth (mix real/synth data) 0.602 (0.566-0.638) 20.1 (14.8-25.3)
@@ 4.6 Use of Pseudo-labelling
The use of synthetic data for semi-supervised pseudo-label training is also explored through the use of a large collection of MPRAGE volumes from the PLORAS study (Price et al., 2010). Results in Figure 7 and Tables 8, 9 and 10 demonstrate an overall positive effect of pseudo labels in the Synth model, with the ARC ensemble result seeing a statistically significant increase in Dice as a result of the new data. The in-domain ATLAS dataset also shows a notable improvement for both models as a result of pseudo labels, with a more marked increase in the baseline model.
@@
Figure 7: Dice and 95th-percentile Hausdorff Distance for all reported datasets, for models trained with/without an additional dataset of MPRAGE images and pseudo-labels.
Table 8: Median results with and without pseudo-label training on the ATLAS test set (N=131). Best score shown in bold. Student’s t distribution 95% confidence intervals given in brackets.
Modality Model Dice HD95
T1w Baseline 0.575 (0.523-0.627) 23.5 (13.9-33.2)
Baseline+Pseudo 0.598 (0.548-0.647) 21.5 (13.8-29.1)
Synth 0.482 (0.431-0.534) 22.6 (12.0-33.3)
Synth+Pseudo 0.496 (0.445-0.547) 22.8 (11.5-34.1)
Table 9: Median results with and without pseudo-label training on the ISLES 2015 test set (N=28). Best score shown in bold. Student’s t distribution 95% confidence intervals given in brackets.
Modality Model Dice HD95
T1w Baseline 0.000 (0.000-0.074) 67.0 (39.7-94.2)
Baseline+Pseudo 0.003 (0.000-0.091) 74.2 (48.3-100.0)
Synth 0.110 (0.012-0.208) 52.6 (30.0-75.1)
Synth+Pseudo 0.179 (0.076-0.283) 44.8 (20.9-68.8)
T2w Baseline 0.000 (0.000-0.005) 62.7 (56.3-69.1)
Baseline+Pseudo 0.001 (0.000-0.007) 64.9 (58.3-71.5)
Synth 0.111 (0.007-0.216) 75.0 (68.4-81.7)
Synth+Pseudo 0.181 (0.071-0.291) 68.3 (61.1-75.4)
FLAIR Baseline 0.000 (0.000-0.000) 65.8 (56.9-74.7)
Baseline+Pseudo 0.000 (0.000-0.003) 73.3 (65.4-81.3)
Synth 0.212 (0.085-0.340) 56.1 (38.7-73.4)
Synth+Pseudo 0.042 (0.000-0.173) 52.9 (30.4-75.4)
DWI Baseline 0.000 (0.000-0.000) 73.1 (57.5-88.7)
Baseline+Pseudo 0.000 (0.000-0.004) 71.6 (67.5-75.7)
Synth 0.056 (0.000-0.144) 84.9 (77.2-92.6)
Synth+Pseudo 0.056 (0.000-0.147) 80.4 (71.6-89.2)
Ensemble Baseline 0.000 (0.000-0.000) 256.0 (237.2-256.0)
Baseline+Pseudo 0.000 (0.000-0.000) 256.0 (231.1-256.0)
Synth 0.423 (0.302-0.545) 47.3 (24.2-70.4)
Synth+Pseudo 0.501 (0.376-0.626) 55.3 (43.7-67.0)
Table 10: Median results with and without pseudo-label training on the ARC test set (N=229). Best score shown in bold. Student’s t distribution 95% confidence intervals given in brackets.
Modality Model Dice HD95
T1w Baseline 0.752 (0.715-0.790) 8.6 (2.3-14.9)
Baseline+Pseudo 0.759 (0.721-0.797) 9.3 (3.8-14.7)
Synth 0.723 (0.684-0.762) 11.0 (4.0-18.0)
Synth+Pseudo 0.743 (0.704-0.782) 10.0 (3.3-16.7)
T2w Baseline 0.004 (0.000-0.014) 74.2 (71.1-77.2)
Baseline+Pseudo 0.014 (0.003-0.024) 72.1 (69.0-75.1)
Synth 0.268 (0.233-0.302) 64.0 (60.5-67.4)
Synth+Pseudo 0.372 (0.332-0.411) 63.1 (59.2-66.9)
FLAIR Baseline 0.120 (0.076-0.163) 63.0 (54.1-71.9)
Baseline+Pseudo 0.145 (0.101-0.189) 70.6 (62.4-78.8)
Synth 0.141 (0.097-0.185) 52.2 (43.3-61.1)
Synth+Pseudo 0.129 (0.084-0.173) 62.1 (53.5-70.7)
Ensemble Baseline 0.034 (0.013-0.054) 40.4 (33.5-47.3)
Baseline+Pseudo 0.124 (0.097-0.150) 33.7 (28.8-38.6)
Synth 0.602 (0.566-0.638) 20.1 (14.8-25.3)
Synth+Pseudo 0.694 (0.658-0.730) 14.4 (9.9-18.9)
@@ 5 Conclusion
In this study, we introduced a novel synthetic data generation and training framework for stroke lesion segmentation, building on the success of prior works in healthy brain parcellation. A model trained using this novel framework was evaluated on a wide range of datasets covering research and clinical data, and chronic and acute stroke pathologies. Results demonstrated that the proposed model is on par with the state-of-the-art for in-domain MPRAGE data, whilst being able to significantly outperform the state-of-the-art model in a variety of out-of-domain scenarios. This was further supported by comparison to a number of state-of-the-art domain adaptation techniques. These domain adaptation techniques consistently led to improvements from the base prediction of our new model, whilst degrading the already poor performance of the MPRAGE-trained model in all scenarios. This provides evidence for the value of our model as a robust ’one size fits all’ base model that can provide an excellent initial prediction with which to adapt or fine-tune to a new domain.
Qualitative investigation also demonstrates the potential of the model for a starting checkpoint in an active-learning framework such as MONAI Label (Diaz-Pinto et al., 2024) where predictions may be refined to generate training data for task-specific fine-tuning. The uncertainty of healthy-tissue segmentations via MC dropout may also provide an effective heuristic for prioritising items to refine/label in such a framework (Nath et al., 2021).
Results also showed the potential benefit of the model for semi-supervised learning through pseudo labelling, where predicted labels may be fed back into the model with much less caution than is typically required when training with real images (Pham et al., 2021; Xu et al., 2024). A simple solution to further reinforce this may be to train a real/fake discriminator to perform automated quality-control and threshold optimisation for the generated binary maps.
@@
LC is supported by the EPSRC-funded UCL Centre for Doctoral Training in Intelligent, Integrated Imaging in Healthcare (i4health) (EP/S021930/1), and the Wellcome Trust (203147/Z/16/Z and 205103/Z/16/Z). IP supported by the Alzheimer’s Association grant number SG-20-678486-GAAIN2. This research was supported by NVIDIA and utilized NVIDIA RTX A6000 48GB.
@@
Appendix A Further segmentation metrics

### v3

@@ 4.6 Ablation: To mix or not to mix?
For all experiments shown thus far, the Synth model used a mix of both synthetic data and the real ATLAS dataset. In order to evaluate whether this decision is justified, an ablation is performed where the Synth model using mixed real/synthetic data is compared to a model trained with only synthetic data. This model is trained in the exact same manner as described previously for the baseline and Synth models. Results for the test datasets ATLAS, ARC and ISLES 2015 are shown in Tables 6, 7 and 8 respectively.
The results reveal a nuanced picture. For in-domain and near-domain T1w data (ATLAS, ARC), mixing with real data provides clear benefits, with pure synthetic training achieving only 19.7% and 46.7% Dice respectively compared to 48.2% and 72.3% for mixed training. However, for several out-of-domain scenarios, pure synthetic training surprisingly often outperforms mixed training: ARC T2w (62.6% vs 26.8%), ISLES T1w (30.4% vs 11.0%), and ISLES FLAIR (37.2% vs 21.2%). This suggests that including real T1w data may inadvertently bias the model toward T1w-specific features, reducing generalisation to other modalities. The T2w and FLAIR improvements with pure synthetic training may reflect these sequences’ different tissue contrast patterns, which are better captured by unbiased synthetic variation than by a model partially trained on T1w-specific features. Further improving the realism of the generated lesions with methods proposed in Liu et al. (2025) may aid in closing the gap for the model trained only on synthetic data, potentially achieving both in-domain performance and out-of-domain generalisation.
@@
Figure 5: Dice and Surface Dice metrics for all reported datasets, for models trained on different combinations of real/synthetic data.
@@
Table 6: Median results for ablation and pseudo-label experiments on the ATLAS test set (N=131). Best score shown in bold. Student’s t distribution 95% confidence intervals given in brackets.
Modality Model Dice (%) Surface Dice (%)
T1w Baseline 57.5 (52.3-62.7) 49.4 (44.5-54.3)
Baseline+Pseudo 59.8 (54.8-64.7) 48.6 (44.0-53.3)
Synth (no real data) 19.7 (15.5-23.9) 14.6 (12.3-16.9)
Synth (Ours) 48.2 (43.1-53.4) 38.1 (33.4-42.7)
Synth+Pseudo 49.6 (44.5-54.7) 38.2 (33.8-42.6)
@@
Table 7: Median results for ablation and pseudo-label experiments on the ARC test set (N=229). Best score shown in bold. Student’s t distribution 95% confidence intervals given in brackets.
Modality Model Dice (%) Surface Dice (%)
T1w Baseline 75.2 (71.5-79.0) 41.7 (39.1-44.4)
Baseline+Pseudo 75.9 (72.1-79.7) 38.2 (35.8-40.7)
Synth (no real data) 46.7 (43.0-50.4) 18.2 (16.5-20.0)
Synth (Ours) 72.3 (68.4-76.2) 33.7 (31.2-36.2)
Synth+Pseudo 74.3 (70.4-78.2) 36.8 (34.3-39.3)
T2w Baseline 0.4 (0.0-1.4) 1.2 (0.8-1.7)
Baseline+Pseudo 1.4 (0.3-2.4) 3.0 (2.5-3.4)
Synth (no real data) 62.6 (58.8-66.5) 30.9 (28.7-33.0)
Synth (Ours) 26.8 (23.3-30.2) 12.0 (10.4-13.6)
Synth+Pseudo 37.2 (33.2-41.1) 14.7 (12.7-16.7)
FLAIR Baseline 12.0 (7.6-16.3) 6.3 (4.7-7.8)
Baseline+Pseudo 14.5 (10.1-18.9) 7.0 (5.5-8.5)
Synth (no real data) 9.6 (7.0-12.3) 7.3 (6.1-8.6)
Synth (Ours) 14.1 (9.7-18.5) 6.6 (4.8-8.4)
Synth+Pseudo 12.9 (8.4-17.3) 6.3 (4.4-8.1)
Ensemble Baseline 3.4 (1.3-5.4) 2.0 (0.8-3.2)
Baseline+Pseudo 12.4 (9.7-15.0) 6.5 (5.1-8.0)
Synth (no real data) 59.7 (56.2-63.1) 26.0 (24.2-27.9)
Synth (Ours) 60.2 (56.6-63.8) 26.3 (24.3-28.4)
Synth+Pseudo 69.4 (65.8-73.0) 34.0 (31.7-36.3)
Table 8: Median results for ablation and pseudo-label experiments on the ISLES 2015 test set (N=28). Best score shown in bold. Student’s t distribution 95% confidence intervals given in brackets.
Modality Model Dice (%) Surface Dice (%)
T1w Baseline 0.0 (0.0-7.4) 0.0 (0.0-4.8)
Baseline+Pseudo 0.3 (0.0-9.1) 0.0 (0.0-4.0)
Synth (no real data) 30.4 (21.3-39.6) 9.4 (4.1-14.8)
Synth (Ours) 11.0 (1.2-20.8) 3.4 (0.0-8.5)
Synth+Pseudo 17.9 (7.6-28.3) 8.0 (2.1-13.8)
T2w Baseline 0.0 (0.0-0.5) 0.3 (0.0-1.0)
Baseline+Pseudo 0.1 (0.0-0.7) 0.4 (0.0-1.1)
Synth (no real data) 7.4 (0.0-18.4) 8.1 (3.0-13.1)
Synth (Ours) 11.1 (0.7-21.6) 7.1 (2.7-11.5)
Synth+Pseudo 18.1 (7.1-29.1) 9.5 (4.4-14.5)
FLAIR Baseline 0.0 (0.0-0.0) 0.0 (0.0-0.0)
Baseline+Pseudo 0.0 (0.0-0.3) 0.0 (0.0-0.4)
Synth (no real data) 37.2 (25.5-48.9) 16.0 (10.8-21.1)
Synth (Ours) 21.2 (8.5-34.0) 14.7 (9.0-20.3)
Synth+Pseudo 4.2 (0.0-17.3) 3.9 (0.0-9.8)
DWI Baseline 0.0 (0.0-0.0) 0.0 (0.0-0.0)
Baseline+Pseudo 0.0 (0.0-0.4) 0.0 (0.0-0.0)
Synth (no real data) 8.2 (0.0-16.8) 5.1 (2.0-8.3)
Synth (Ours) 5.6 (0.0-14.4) 4.1 (0.8-7.4)
Synth+Pseudo 5.6 (0.0-14.7) 4.3 (0.8-7.8)
Ensemble Baseline 0.0 (0.0-0.0) 0.0 (0.0-0.0)
Baseline+Pseudo 0.0 (0.0-0.0) 0.0 (0.0-0.0)
Synth (no real data) 27.2 (15.7-38.8) 15.6 (10.3-20.9)
Synth (Ours) 42.3 (30.2-54.5) 19.1 (12.7-25.4)
Synth+Pseudo 50.1 (37.6-62.6) 24.9 (18.7-31.1)
@@ 4.7 Use of Pseudo-labelling
The use of synthetic data for semi-supervised pseudo-label training is also explored through the use of the PLORAS MPRAGE dataset (N=1159), distinct from the PLORAS hospital scans used in the Experiments. This approach leverages a key advantage of our framework: because synthetic training decouples image generation from label accuracy, the model can tolerate and potentially benefit from imperfect pseudo-labels. Results in Figure 6 and Tables 6, 7 and 8 demonstrate an overall positive effect of pseudo labels in the Synth model, with the exception of FLAIR images in ISLES 2015.
Notable improvements include: ARC T2w increasing from 26.8% to 37.2% Dice, ISLES T1w from 11.0% to 17.9%, and ARC ensemble from 60.2% to 69.4%. The baseline model shows minimal improvement with pseudo-labels in out-of-domain scenarios (e.g., ISLES ensemble remains at 0.0%), suggesting that pseudo-labelling cannot overcome fundamental domain shift. The FLAIR degradation in ISLES 2015 (21.2% to 4.2%) may indicate that pseudo-labels from chronic MPRAGE data introduce biases incompatible with acute FLAIR lesion appearance. The in-domain ATLAS dataset also shows a notable improvement for both models as a result of pseudo labels, with a more marked increase in the baseline model, suggesting pseudo-labels effectively expand the training distribution when the domain gap is small.
@@
Figure 6: Dice and Surface Dice metrics for all reported datasets, for models trained with/without an additional training dataset of MPRAGE images and pseudo-labels.
These results across four diverse datasets - from research-quality ATLAS to challenging clinical PLORAS scans - reveal consistent patterns that warrant deeper analysis of the fundamental principles underlying domain robustness.
5 Discussion
Our results reveal fundamental principles about domain robustness in medical image segmentation. The consistent failure of domain adaptation to rescue baseline models - even with oracle selection of the optimal method - demonstrates that post-hoc adaptation cannot substitute for domain-invariant training. This finding challenges prevailing assumptions about the sufficiency of test-time adaptation for clinical deployment.
The compound effect observed with Synth+DAE suggests that synthetic pre-training and test-time adaptation address different aspects of domain shift. Synthetic training provides feature representations invariant to appearance changes, while DA methods fine-tune these representations to specific test-time distributions. However, this synergy only emerges when the base model already possesses sufficient robustness - when initial predictions fail catastrophically (as with baseline models on ISLES 2015 and PLORAS), no amount of adaptation can recover performance.
The variable benefit of multi-modal ensembling - substantial for ISLES 2015 but mixed for ARC - indicates that fusion strategies must consider dataset-specific characteristics rather than assuming universal improvement. Clinical deployment should balance single-modality robustness with opportunistic multi-modal fusion.
6 Limitations and Future Directions
Appendix A confirms that 20–40% of lesions in several cohorts are multimodal, underscoring the central limitation of our single-Gaussian sampling. While our spatially-varying approach introduces intra-lesion heterogeneity, it does not address the multi-modal nature of stroke appearance. Future implementations should explore mixture models to capture lesions that simultaneously exhibit hyper- and hypo-intense regions within the same pathology. It may also be advantageous to paste several lesions per subject, each assigned its own independently sampled intensity profile. This would allow a single synthetic brain to display chronic hypointense scars alongside acute hyperintense infarcts, better reflecting the heterogeneous lesion chronology typically observed in stroke cohorts.
Additional qualitative results in Appendix Figures 18 - 23 display further positive and negative results across the available image modalities within the PLORAS dataset. A number of failure modes may be observed here, such as missing cerebellar lesions or under-segmenting large hemispheric lesions.
@@ 7 Conclusion
In this study, we introduced a novel synthetic data generation and training framework for stroke lesion segmentation, building on the success of prior works in healthy brain parcellation. A model trained using this novel framework was evaluated on a wide range of datasets covering research and clinical data, and chronic and acute stroke pathologies. Our experiments demonstrate that synthetic pre-training provides fundamental robustness unachievable through test-time adaptation alone. While accepting a 9.3% median Dice reduction in-domain (48.2% vs 57.5%), our approach maintains performance where conventional methods fail entirely. Even with oracle knowledge of the optimal domain adaptation method - an unrealistic scenario in practice - conventionally-trained models cannot match our synthetic approach in out-of-domain settings.
Our results demonstrate that even imperfect appearance modelling can provide substantial benefits for cross-modality segmentation. Incorporating recent contemporary work on realistic lesion simulation (Liu et al., 2025) may further improve performance both in- and out-of-distribution.
Qualitative results in Figure 4 demonstrate the potential of the model as a starting checkpoint in an active-learning framework such as MONAI Label (Diaz-Pinto et al., 2024), where predictions may be refined to generate training data for task-specific fine-tuning. The uncertainty of healthy-tissue segmentations via MC dropout may also provide an effective heuristic for prioritising items to refine/label in such a framework (Nath et al., 2021).
Results also showed the potential benefit of the model for semi-supervised learning through pseudo labelling, where predicted labels may be fed back into the model with much less caution than is typically required when training with real images (Pham et al., 2021; Xu et al., 2024). The ablation study revealed that pure synthetic training can outperform mixed training for certain out-of-domain modalities, suggesting opportunities for further optimisation of the real/synthetic data balance. A simple solution to further reinforce this may be to train a real/fake discriminator to perform automated quality-control and threshold optimisation for the generated binary maps.
Our approach reduces reliance on domain-specific training data and helps bridge the gap between research-grade and clinical scans to improve clinical stroke neuroimaging workflows, providing a foundation for more robust and widely applicable lesion segmentation tools.
@@
LC is supported by the EPSRC-funded UCL Centre for Doctoral Training in Intelligent, Integrated Imaging in Healthcare (i4health) (EP/S021930/1), and the Wellcome Trust (203147/Z/16/Z and 205103/Z/16/Z). IP supported by the Alzheimer’s Association grant number SG-20-678486-GAAIN2. CP is funded by Wellcome (203147/Z/16/Z, 205103/Z/16/Z and 224562/Z/21/Z). This research was supported by NVIDIA and utilised NVIDIA RTX A6000 48GB.
@@
Appendix A Lesion-intensity unimodality
Motivation. Most classical stroke-segmentation pipelines, including our synthetic generator, implicitly assume that the grey-level distribution of a lesion is unimodal. That assumption underpins common appearance priors such as single-Gaussian modelling and simple intensity normalisation. However, chronic infarcts often contain a mixture of tissue constituents, and in CT the same infarct can include both hypo- and hyper-dense cores. Quantifying how often real lesions violate unimodality therefore informs whether richer mixture-based priors are warranted.
Experiment. For every connected component in each dataset, we extracted the raw voxel intensities and applied Hartigan’s dip test (Hartigan and Hartigan, 1985) to the empirical one-dimensional distribution. A lesion was labelled unimodal when the null hypothesis of unimodality could not be rejected at $\alpha=0.05$; otherwise it was classified as multimodal. For efficiency, we randomly subsampled at most 2048 voxels per lesion before testing. Table 9 summarises the counts.
Findings. Across MRI datasets the majority of lesions were indeed unimodal, but a non-negligible tail of multimodal cases emerged:
  • ATLAS (chronic MPRAGE T1w) showed $95\%$ unimodal lesions; only 80 of 1818 connected components exhibited multimodality, in line with mature cavities whose signal is dominated by a single CSF-like class.
  • ISLES 2015 exhibited 12-38% multimodal lesions depending on modality, reflecting mixed-phase tissue in a sub-acute cohort.
  • ARC (chronic research scans) retained a high unimodal rate in T1w/T2w but FLAIR contained 39% multimodal lesions.
  • PLORAS revealed the strongest departure: while T2w/FLAIR remained largely unimodal, CT lesions were predominantly multimodal (58%), underscoring how density heterogeneity dominates in CT.
Implications. A single-Gaussian appearance prior is largely adequate for T1w/T2w MRI, where $\geq 80\%$ of lesions were unimodal; it begins to break down for modalities that emphasise subtle tissue heterogeneity - most notably FLAIR and especially CT. Adopting explicit mixture-based priors, or synthetically sampling lesions with mixed intensities, therefore represents a principled next step for improving model robustness across modalities and imaging protocols.
Table 9: Counts of connected‐component lesions that are unimodal or multimodal in each dataset–modality pair. For every lesion larger than 10 voxels we sampled at most 2048 raw intensity values and ran Hartigan’s dip test; failure to reject the null hypothesis of unimodality at $\alpha=0.05$ yields the Unimodal label, otherwise Multimodal. ATLAS is split into the subjects used for model development (T1w [Train]) and the held-out evaluation set (T1w [Validation]).
Dataset Modality Unimodal Multimodal
ATLAS T1w [Train] 1384 65
T1w [Validation] 354 15
ISLES2015 T1w 52 9
T2w 46 15
FLAIR 54 7
DWI 38 23
ARC T1w 286 64
T2w 335 48
FLAIR 90 57
PLORAS T2w 203 14
FLAIR 712 29
CT 248 342
Appendix B Further segmentation metrics
Additional metrics are reported below in Appendix Tables 10 - 13 for all datasets. Here we report individual scores for each DA method, for both the baseline model and our Synth model.